A generator with a circular coil of 20 turns of area 2.0 x 10-2 m2
is immersed in
a 0.40 T magnetic field and rotated with a frequency of 50 Hz. Find (i) The Flux
linkage (ii) The EMF (iii) the maximum EMF which is produced during a cycle
1
Expert's answer
2020-08-12T15:26:48-0400

As per the question,

Number of turns (n)=20

Area "(A) =2.0 \times 10^{-2} m^2"

Magnetic field "(B)=0.4T"

Rotational frequency "(f)=50Hz"

(i) The flux linkage "(\phi) =nBA"

Now substituting values in the above,

"\Rightarrow \phi=20\times 2.0\times 10^{-2}\times 0.4 \ \ Tm^2"

"=16.0\times 10^{-2}Tm^2"

(ii) Emf produced by the generator "e=nBA\omega"

"e=nBA\times 2\pi f"

Now, substituting the values,

"\Rightarrow e=20\times 0.4\times 2.0\times 10^{-2}\times 2\times 3.14\times 50 Tm^2 rad/sec"

"\Rightarrow e=50.24V"